# Discrete-Time CRLB-based Power Allocation for CF MIMO-ISAC with Joint Localization and Velocity Sensing

### v2

@@ Abstract
In this paper, we investigate integrated sensing and communication (ISAC) in a cell-free (CF) multiple-input multiple-output (MIMO) network, where each access point functions either as an ISAC transmitter or as a sensing receiver. We devote into the ISAC sensing metric using the discrete-time signal-based Cramér–Rao lower bounds (CRLBs) for joint location and velocity estimation under arbitrary power allocation ratios under the deterministic radar cross section assumption (RCS). Then, we consider the power allocation optimization problem for the CF MIMO-ISAC as the maximization of the communication signal-to-interference-plus-noise ratio (SINR), subject to CRLB-based sensing constraints and per-transmitter power limits. To solve the resulting nonlinear and non-convex problem, we propose a penalty function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS), and an alternative method based on a modified steepest descent approach (PP-MSD-ILS). We show that the proposed algorithms are scalable and can be extended to a broad class of optimization problems involving nonlinear inequality constraints and affine equality constraints. In addition, we extend the PP-MCG-ILS algorithm to the pure sensing scenario, where a penalty function-based normalized conjugate gradient algorithm (P-NCG-ILS) is developed for sensing power minimization. Finally, we analyze the convergence behavior and qualitatively compare the computational complexity of the proposed algorithms. Simulation results confirm the accuracy of the derived CRLBs and demonstrate the effectiveness of the proposed power allocation strategies in enhancing both sensing and overall ISAC performance.
@@ I Introduction
However, most existing CRLB formulations are built upon continuous-time signal models, whereas practical radar systems operate in the discrete-time domain, where the sampling rate critically affects sensing accuracy. Our prior work [30] addresses this gap by introducing a discrete-time likelihood function over Gaussian channels, offering a more realistic modeling framework. Building on this foundation, this paper devotes into the CRLBs derivation for joint location and velocity estimation under the deterministic RCS assumption, covering both single- and multi-target cases in CF MIMO-ISAC systems.
@@
These considerations motivate us to consider a more comprehensive and realistic setting: joint location and velocity estimation in CF MIMO-ISAC systems under a discrete-time signal model with a deterministic RCS assumption. The CRLBs are serve as effective sensing metrics to guide PA across distributed APs, with the goal of maximizing communication SINR. This leads to a class of non-convex and highly nonlinear optimization problems. To address this issue, we develop a tailored optimization framework that integrates penalty methods, projection operations, and gradient-based search, enabling efficient and reliable algorithmic solutions under complex practical constraints. The main contributions are as follows:
@@
  • We derive novel closed-form approximate CRLBs for joint location and velocity estimation under arbitrary PA strategies. These CRLBs are rigorously validated against maximum likelihood estimation (MLE) results, confirming their effectiveness as practical sensing performance metrics. Based on these metrics, we formulate the PA optimization problems for ISAC, aiming at communication SINR maximization while ensuring joint CRLB and power budget constraints.
@@
  • To tackle the non-convexity and nonlinearity in PA optimization for ISAC, we propose a penalty-function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS) together with PP-MSD-ILS algorithm based on the modified steepest descent algorithm (MSD). We further propose a penalty-function-based normalized modified conjugate gradient algorithm with inexact line search (P-NCG-ILS), which can effectively solve the power minimization problem subject to sensing accuracy constraints.
  • We provide a comprehensive analysis of their convergence behavior and computational complexity. The design ensures stable updates, rapid convergence and low computational overhead. Additionally, we show that the proposed PP-MCG-ILS and P-NCG-ILS algorithms can be broadly scalable to a class of constrained non-convex optimization problems involving affine equality constraints (handled via projection) and nonlinear inequality constraints (handled via penalties).
  • Extensive simulations validate the performance of the developed algorithms in both ISAC and sensing-only scenarios. The PP-MCG-ILS algorithm consistently exhibits fast convergence, robust numerical stability, and optimal PA efficiency across various conditions, as compared to benchmark methods. Results also highlight important trade-offs between sensing and communication performance and confirm that the proposed methods can meet stringent sensing QoS while maximizing communication throughput or minimizing power usage.
@@
Section II introduces the received signal model for the CF MIMO-ISAC system. Section III-B formulates the power allocation problem for CF MIMO-ISAC. Section IV presents the proposed power allocation algorithms for both CF MIMO-ISAC and the sensing-only scenario. Section VI analyzes the convergence properties and computational complexity of the proposed algorithms. Section VII provides simulation results to validate the theoretical analysis and algorithmic performance. Finally, Section VIII concludes the paper.
[…]
@@ II Signal Model
The Doppler frequency induced by the relative radial velocity of the target with respect to transmitter $n$ and receiver $k$ is computed as
where $\lambda$ denotes the wavelength. Denote $\tau_{n,k}$ as the propagation delay between the $n$th transmitter and the $k$th receiver, as reflected by the target. Then, we have
@@
where $\tau_{n}$ and $\tau_{k}$ represent the propagation delays from the $n$th transmitter to the $q$th target, and from the target to the $k$th receiver, respectively, and $c$ denotes the speed of light.
[…]
@@
where $P$ denotes the total available energy, $\rho_{n}$ is the PA factor with $\sum_{n=1}^{N}\rho_{n}=1$, ${b}_{n}\in\mathbb{C}$ is the normalized transmit beamforming weight with $|{b}_{n}|^{2}=1$, and $\varsigma\in\mathbb{C}$ has unit variance $\mathbb{E}\{|\varsigma|^{2}\}=1$. Let ${\bm{\rho}}=[\rho_{1},\rho_{2},\cdots,\rho_{N}]^{\rm T}$ be the vector of the PA coefficient of the $N$ transmitters.
Sensing model: For target-reflected reception, we refer the term “RCS” as encompassing both channel fading and radar RCS. Denote the RCS associated with transmitter $n$ and receiver $k$ as ${\alpha}_{n,k}$. We consider the Swerling-I sensing model [37], i.e. the deterministic RCS assumption where RCSs are assumed to be deterministic but unknown during the observation time [29, 28]. Then, the lowpass equivalent of the reflected signal from the $n$th transmitter by a target received at receiver $k$ is given by [^1] The transmitted signals over different transmitters are orthogonal and maintain approximate orthogonality over different delays and Doppler frequencies such that the signals contributed from different transmitters can be separated at each receiver [29, 26].
@@
where $g_{n}$ is the channel fading between the transmitter $n$ and the user, and $z(t)$ denotes the interference-plus-noise with zero mean and variance $\sigma_{z}^{2}$. Then, the SINR at the user terminal can be given by
@@
where $\delta=\frac{P}{T_{\rm eff}\sigma_{z}^{2}}$ is the total SINR and ${\bm{g}}$ is the channel gain [^2] The communication channels are assumed to be estimated a priori using pilot signals [24, 35, 34, 36]. with its $n$th element being $|g_{n}|^{2}$. Note that $(\mathrm{i})$ holds when using a normalized conjugate beamforming with ${b}_{n}={g}^{*}_{n}/|g_{n}|$.
@@ III The Proposed Sensing Metric and Problem Formulation
In this section, we present an approximate CRLB for CF MIMO radar under arbitrary power allocations.
@@
[…]
@@ III-A Sensing Metric
Unlike the sensing performance metrics employed in [24, 25, 34, 35, 36], which only consider target detection or location estimation, we consider the joint location and velocity CRLB as the sensing metric, developed based on the discrete-time reception of the reflected signal under the deterministic RCS assumption. In particular, we have derived the approximate CRLBs for both location and velocity in [30] for simplified sensing metrics. They are respectively given by
@@
The detailed expression of ${\bm{P}}$ and ${\bm{V}}$ are given at the top of the page, respectively, and
@@
The relevant waveform-related parameters $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, and geometric-spreading-related factors $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, are detailed in Appendix A.
As observed from (61), the PA coefficients play a pivotal role in the formulation of the CRLBs. Meanwhile, these coefficients directly affect communication SINR, as described in (7). Unfortunately, the PA optimization relevant to the CRLB is intractable due to the implicit and complex forms of the expressions in (11) and (12). To facilitate analysis and optimization, we reformulate the CRLB expressions into a more explicit and tractable form.
We denote the $(i,j)$th elements of matrices ${\bm{P}}$, ${\bm{V}}$ and ${\bm{Y}}$ as $p_{ij}$, $v_{ij}$ and $y_{ij}$, respectively. Then, one can simplify them as
@@
where $w_{n}^{p_{ij}}$, $w_{n}^{v_{ij}}$ and $w_{n}^{y_{ij}}$ are the $n$th element of the weight vector ${\bm{w}}_{p_{ij}}$, ${\bm{w}}_{v_{ij}}$ and ${\bm{w}}_{y_{ij}}$, respectively. These weights can be readily obtained from the definitions of ${\bm{P}}$, ${\bm{V}}$ and ${\bm{Y}}$, by extracting their PA coefficient $\rho_{n}$. For example, from (13), we have $w_{n}^{p_{11}}$ as
@@
where $|{\bm{U}}|=u_{11}u_{22}-u_{12}u_{21}$. The relevant elements $u_{ij}$ are given respectively by
$\small u_{11}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (17)
@@
Note that the newly derived CRLB expressions in (16) and (19) represent explicit functions of the PA vector, thereby serving as effective and tractable sensing performance metrics.
@@ III-B Problem Formulation
Let ${\bm{\rho}}_{\min}=[\rho_{\min,1},\rho_{\min,2},\cdots,\rho_{\min,N}]^{\rm T}$ denote the minimum PA coefficient to achieve the sufficiently large SNR for tight CRLBs [34, 36], and ${\bm{\rho}}_{\max}=[\rho_{\max,1},\rho_{\max,2},\cdots,\rho_{\max,N}]^{\rm T}$ denote the maximum PA coefficient due to the power budgets of individual transmitters [38, 36]. In this paper, we consider the problem of maximizing the communication SINR, subject to constraints on the sensing performance, i.e.,
@@
where $\delta_{l}^{2}$ and $\delta_{v}^{2}$ denote the corresponding location and velocity CRLB thresholds.
@@ IV Power Allocation Algorithm for CF MIMO-ISAC
In this section, we investigate the optimal PA strategy for maximizing the ISAC performance, as formulated in (22a)-(22e). Unfortunately, the strong nonlinearity and non-convexity of the CRLB-based constraints in (22d) and (22e) render the problem analytically intractable. To overcome this challenge, we adopt a CG iterative algorithm, enhanced with a penalty function and projection operator, to provide an efficient numerical solution.
To derive a numerical solution, we start from converting the original constrained optimization problem into an unconstrained one. To this end, we first introduce a penalty (function) for violating the nonlinear inequality constraints in (22c)-(22e), defined as [39]

### v3

@@ Abstract
This paper investigates integrated sensing and communication (ISAC) in a cell-free (CF) multiple-input multiple-output (MIMO) network, in which each access point operates either as an ISAC transmitter or as a sensing receiver. Discrete-time signal-based Cramér–Rao lower bounds (CRLBs) are employed as sensing performance metrics for joint location and velocity estimation under arbitrary power allocation ratios, assuming a deterministic radar cross section (RCS) model. Based on these metrics, a power allocation problem for CF MIMO-ISAC is formulated to maximize the communication signal-to-interference-plus-noise ratio (SINR), subject to CRLB-based sensing constraints and per-transmitter power limits. To address the resulting nonlinear and nonconvex optimization problem, a penalty-function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS) is developed, together with an alternative method based on modified steepest descent (PP-MSD-ILS). The proposed algorithms are applicable to a class of constrained optimization problems involving nonlinear inequality constraints and affine equality constraints. In addition, the PP-MCG-ILS algorithm is extended to a sensing-only scenario, where a penalty-function-based normalized conjugate gradient algorithm (P-NCG-ILS) is employed for sensing power minimization. The convergence behavior of the proposed algorithms is analyzed, and their computational complexity is qualitatively evaluated. Simulation results validate the accuracy of the derived CRLBs and demonstrate the effectiveness of the proposed power allocation strategies in enhancing sensing performance and overall ISAC performance.
@@ I Introduction
However, most existing CRLB formulations are built upon continuous-time signal models, whereas practical radar systems operate in the discrete-time domain, where the sampling rate critically affects sensing accuracy. Our prior work [40] addresses this gap by introducing a discrete-time likelihood function over Gaussian channels, offering a more realistic modeling framework. Building on this foundation, this paper devotes into the CRLBs derivation for joint location and velocity estimation under the deterministic RCS assumption in CF MIMO-ISAC systems.
@@
These considerations motivate us to consider a more comprehensive and realistic setting: joint location and velocity estimation in CF MIMO-ISAC systems under a discrete-time signal model with a deterministic RCS assumption. The CRLBs are served as effective sensing metrics to guide PA across distributed APs, with the goal of maximizing communication SINR. This leads to a class of non-convex and highly nonlinear optimization problems. To address this issue, we develop a tailored optimization framework that integrates penalty methods, projection operations, and gradient-based search, enabling efficient and reliable algorithmic solutions under complex practical constraints. The main contributions of this work are summarized as follows:
@@
  • We develop tractable closed-form approximations of the CRLB for joint location and velocity estimation under arbitrary power allocation (PA) strategies. The accuracy of the proposed approximations is validated through comparisons with maximum-likelihood estimation (MLE) results, indicating their suitability as sensing performance metrics for ISAC-oriented resource allocation. Based on these metrics, PA optimization problems are formulated to maximize communication SINR subject to sensing accuracy and power budget constraints.
@@
  • To address the nonconvex and nonlinear nature of the resulting PA optimization problems, we introduce a penalty-function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS), along with a PP-MSD-ILS algorithm based on modified steepest descent. In addition, a penalty-function-based normalized conjugate gradient algorithm with inexact line search (P-NCG-ILS) is developed for power minimization in sensing-only scenarios under prescribed sensing accuracy constraints.
  • We provide a detailed analysis of the convergence behavior and computational complexity of the proposed algorithms. Under standard smoothness assumptions, the analysis establishes well-defined step size selection, stable iterative updates, and efficient numerical behavior. The algorithmic framework also admits straightforward extension to related constrained nonconvex optimization problems.
  • Simulation results are provided to evaluate the proposed algorithms in both ISAC and sensing-only scenarios. The results demonstrate that the proposed methods achieve reliable convergence and competitive PA performance relative to representative benchmark schemes, while revealing the inherent trade-offs between sensing accuracy and communication performance.
@@
Section II introduces the received signal model for the CF MIMO-ISAC system. Section III formulates the PA problem for CF MIMO-ISAC based on the proposed metrics. Section IV presents the proposed PA algorithms for both CF MIMO-ISAC and the sensing-only scenario. Section VI analyzes the convergence properties and computational complexity of the proposed algorithms. Section VII provides simulation results to validate the theoretical analysis and algorithmic performance. Finally, Section VIII concludes the paper.
Notation: $\mathbb{R}$ and $\mathbb{C}$ denote the field of real and complex numbers, respectively. Scalars are denoted by lower-case letters, vectors and matrices respectively by lower- and upper-case boldface letters. The conjugate and transpose are denoted by $(\cdot)^{*}$ and $(\cdot)^{\rm T}$, respectively. $\mathbb{E}\{\cdot\}$ and $|\cdot|$ denote the mathematical expectation and element-wise modulus, respectively. $\|\cdot\|$ denotes the $l_{2}$ norm of a matrix. The notations $\min\{\cdot\}$ and $\max\{\cdot\}$ denote the minimum and maximum element of the enclosed set $\{\cdot\}$, respectively.
@@ II Signal Model
Denote $\tau_{n,k}$ as the propagation delay between the $n$th transmitter and the $k$th receiver, as reflected by the target. Then, we have
@@
where $\tau_{n}$ and $\tau_{k}$ represent the propagation delays from the $n$th transmitter to the $q$th target, and from the target to the $k$th receiver, respectively, and $c$ denotes the speed of light. The Doppler frequency induced by the relative radial velocity of the target with respect to transmitter $n$ and receiver $k$ is
where $\lambda$ denotes the wavelength.
A waveform $s_{n}(t)$, with unit energy, i.e., $\varepsilon_{s}=\int|{s}_{n}(t)|^{2}\,dt=1$, is considered for both sensing and communication at the $n$th transmitter. Then, the baseband ISAC transmitted signal at transmitter $n$ is denoted as
@@
where $P$ denotes the total available energy, $\rho_{n}$ is the PA factor with $\sum_{n=1}^{N}\rho_{n}=1$, ${b}_{n}\in\mathbb{C}$ is the normalized precoding weight with $|{b}_{n}|^{2}=1$, and $\varsigma\in\mathbb{C}$ has unit variance $\mathbb{E}\{|\varsigma|^{2}\}=1$. Let ${\bm{\rho}}=[\rho_{1},\rho_{2},\cdots,\rho_{N}]^{\rm T}$ be the PA coefficient vector over the $N$ transmitters.
Sensing model: For target-reflected reception, we hereafter refer to the term “RCS” as encompassing both channel fading and radar RCS. Denote the RCS associated with transmitter $n$ and receiver $k$ as ${\alpha}_{n,k}$. The sensing model under the deterministic RCS assumption is considered where RCSs are assumed to be deterministic but unknown during the observation time [3, 16, 38]. Then, the lowpass equivalent of the reflected signal from the $n$th transmitter by a target received at receiver $k$ is given by [^1] The transmitted signals over different transmitters are orthogonal and maintain approximate orthogonality over different delays and Doppler frequencies such that the signals contributed from different transmitters can be separated at each receiver [3, 19].
@@
where $g_{n}$ is the channel fading [^2] The communication channels are assumed to be estimated a priori using pilot signals [9, 22, 1, 2]. between the transmitter $n$ and the user, and $z(t)$ denotes the interference-plus-noise with zero mean and variance $\sigma_{z}^{2}$. Then, the SINR at the user terminal can be given by
@@
where $\delta=\frac{P}{T_{\rm eff}\sigma_{z}^{2}}$ is the total loss-free SINR and $\bm{g}$} is the squared channel fading with its $n$th element being $|g_{n}|^{2}$. Note that $(\mathrm{i})$ holds when using a normalized conjugate precoding with ${b}_{n}={g}^{*}_{n}/|g_{n}|$.
Because both sensing and communication channels are assumed quasi-static/deterministic within the coherence observation time, the deterministic RCS model and the estimated communication gains remain valid for each optimization cycle.
@@ III The Proposed Sensing Metric and Problem Formulation
In this section, we propose a sensing metric of the joint location and velocity estimation based on our derived approximate CRLB for CF MIMO radar under arbitrary PAs.
@@
$\small\bar{f_{n}^{2}}=\frac{1}{4\pi^{2}}\int\bigg|\frac{d{s_{n}(t-\tau_{n,k})}}{d{\tau_{n,k}}}\bigg|^{2}\,dt=\int f^{2}|S_{n}(f)|^{2}\,df,$ (8)
and $\bar{t_{n}^{2}}=\int t^{2}|s_{n}(t)|^{2}\,dt$, respectively. The cross-term between time and frequency is further defined as
$\small\sigma_{tf}=\int{ts_{n}^{*}(t-\tau_{n,k})}\frac{d{s_{n}(t-\tau_{n,k})}}{d{\tau_{n,k}}}\,dt.$ (9)
$|s_{n}(t)|^{2}$ and $|S_{n}(f)|^{2}$ can be interpreted as the probability density functions for consecutive random variables $t$ and $f$, respectively, since $\int|S_{n}(f)|^{2}\ df=\int|s_{n}(t)|^{2}\ dt=1$. Based on this interpretation, we define the average time and average frequency of the transmission waveform as
$\small\bar{f}_{n}=\int f|S_{n}(f)|^{2}\ df\quad\text{and}\quad\bar{t}_{n}=\int t|s_{n}(t)|^{2}\ dt,$ (10)
@@ III-A Sensing Metric
Unlike the sensing performance metrics employed in [9, 13, 1, 22, 2], which only consider target detection or location estimation, we consider the joint location and velocity CRLB as the sensing metric, developed based on the discrete-time reception of the reflected signal under the deterministic RCS assumption. We derive the approximate CRLBs of the joint location and velocity estimation [^3] For completeness, the detailed derivations are referred to [40]., respectively, given by
@@
where $\bm{P}$} and $\bm{V}$} are presented at the top of the page, respectively, and
@@
The waveform-related parameters $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, and geometric-spreading-related factors $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, are detailed in Appendix A.
As described in (7), the PA coefficients directly affect communication SINR. However, the PA optimization relevant to the CRLB is intractable due to the implicit and complex forms of the expressions in (11) and (12). To facilitate analysis and optimization, we reformulate the CRLB expressions into a more explicit and tractable form regarding the PA coefficients.
We denote the $(i,j)$th elements of matrices $\bm{P}$}, $\bm{V}$} and $\bm{Y}$} by $p_{ij}$, $v_{ij}$ and $y_{ij}$, respectively. They can be rewritten as
@@
where $w_{n}^{p_{ij}}$, $w_{n}^{v_{ij}}$ and $w_{n}^{y_{ij}}$ are the $n$th element of the weight vector ${\bm{w}}_{p_{ij}}$, ${\bm{w}}_{v_{ij}}$ and ${\bm{w}}_{y_{ij}}$, respectively. These weights can be readily obtained by extracting $\rho_{n}$ from the element-wise definitions of $\bm{P}$}, $\bm{V}$} and $\bm{Y}$}. For example, we have
@@
where $|{\bm{U}}|=u_{11}u_{22}-u_{12}u_{21}$. The elements $u_{ij}$ are given by
$u_{11}=\frac{1}{|{\bm{Y}}|}\big[{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big],$ (17)
@@
The newly derived CRLB expressions in (16) and (19) represent explicit functions of the PA coefficient vector $\bm{\rho}$}, thereby serving as effective and tractable sensing performance metrics.
@@ III-B Problem Formulation
This paper considers the problem of maximizing the communication SINR in (7), subject to constraints on the performance of joint location and velocity estimation, i.e.,
@@
where $\delta_{l}^{2}$ and $\delta_{v}^{2}$ denote the corresponding location and velocity CRLB thresholds  [^4] The thresholds can be specified by referring to the CRLB values under uniform PA..
Note that ${\bm{\rho}}_{\min}=[\rho_{\min,1},\rho_{\min,2},\ldots,\rho_{\min,N}]^{\mathrm{T}}$ denote the minimum PA coefficient vector required to ensure a sufficiently high SNR for achieving tight CRLBs [1, 2], and ${\bm{\rho}}_{\max}=[\rho_{\max,1},\rho_{\max,2},\ldots,\rho_{\max,N}]^{\mathrm{T}}$ denote the maximum PA coefficient vector determined by the individual transmitters’ power budgets [17, 2]. Specifically, (22b) and (22c) represent the total power and per-transmitter power constraints, respectively, while (22d) and (22e) correspond to the location and velocity CRLB constraints, respectively.
@@ IV Power Allocation Algorithm for CF MIMO-ISAC
We now investigate the PA strategy by solving the optimization model formulated in (22a)–(22e). The strong nonlinearity and nonconvexity of the CRLB-based constraints in (22d) and (22e) make the problem analytically intractable. Although numerous numerical optimization methods developed in the literature in past several decades, these methods cannot be directly applied to the nonlinear and nonconvex constrained PA optimization problem considered in this work. To address this challenge, we develop a penalty function and projection operator enhanced conjugate gradient (CG)–based iterative algorithm, to yield an efficient numerical solution.
IV-A Unconstrained Optimization Enabled by Penalty Function and Projection Operator
To facilitate optimization, we start from converting the original constrained problem into an unconstrained one. Firstly, we consider incorporating the nonlinear inequality constraints in (22c)-(22e) onto the target function (22a), by introducing a penalty (function) for violating these constraints. It can be defined as [32]